If you had 0.956 moles of a nonelectrolyte solute in 345 grams of water. What is the new freezing point? (The Kf of water is 1.86 °C/m )
1
Expert's answer
2020-06-05T05:13:41-0400

Formula used in calculating freezing point is

"\Delta Tf=iKfm"

Molarity of the solution

"M=\dfrac{m}{v}"


no.of moles m = 0.956

volume of water = 0.345 kg

"M=\dfrac{0.956}{0.345}" =2.77


Kf = 1.86


Therefore;

"\Delta Tf=1.86*2.77" =5.15


The freezing point is 5.150 C